OpenStudy (anonymous):

can somebody please explain how 16^(-1/2) = 1/4?

OpenStudy (anonymous):

It should equal 4 not 1/4...unless you are looking for the inverse.

OpenStudy (anonymous):

\[16^{ -1/2 }=1/4\] so \[1/16^{1/2}=1/4\] then \[1/\sqrt{16}=1/4\]

OpenStudy (anonymous):

\[a^{ -1 }=\frac{ 1 }{ a }\]

OpenStudy (anonymous):

\[a^{1/2}=\sqrt a\]

OpenStudy (anonymous):

ok great :) thank you again!

OpenStudy (anonymous):

no problem :)